# Risk-Budgeted Control Framework for Balanced Performance and Safety in Autonomous Vehicles

### v1

@@ Abstract
This paper presents a risk–budgeted monitor with a control framework that certifies safety for autonomous driving. In this process, a sliding window is proposed to monitor for insufficient barrier residuals or nonzero tail risk, ensuring system safety. When the safety margin deteriorates, it triggers switching the safety constraint from a performance-based relaxed-control barrier function (R-CBF) to a conservative conditional value at risk (CVaR-CBF) to address the safety concern. This switching is governed by two real–time triggers: Feasibility–Triggered (FT) and Quality–Triggered (QT) conditions. In the FT condition, if the R-CBF constraint becomes infeasible or yields a suboptimal solution, the risk monitor triggers the use of the CVaR constraints for the controller. In the QT condition, the risk monitor observes the safety margin of the R-CBF solution at every step, regardless of feasibility. If it falls below the safety margin, the safety filter switches to the CVaR-CBF constraints.
The proposed framework is evaluated using a model predictive controller (MPC) for autonomous driving in the presence of autonomous vehicle (AV) localization noise and obstacle position uncertainties. Multiple AV-pedestrian interaction scenarios are considered, with 1,500 Monte Carlo runs conducted for all scenarios. In the most challenging setting with pedestrian detection uncertainty of 5m, the proposed framework achieves a 94-96% success rate of not colliding with the pedestrians over 300 trials while maintaining the lowest mean cross–track error (CTE=3.2-3.6 m) to the reference path. The reduced CTE indicates faster trajectory recovery after obstacle avoidance, demonstrating a balance between safety and performance.
@@ I Introduction
Safety, despite uncertainties, is a crucial aspect when developing controllers for Automated vehicles (AVs). These uncertainties drastically increase the computational demand, and the problem becomes challenging when the controller feedback frequency typically needs to operate between 20205050 Hz [1].
Control Barrier Functions (CBFs) in quadratic programming (QP) formulation is one of the most fundamental and widely used strategies in the real-time implementation of obstacle-avoidance, safety-guaranteeing controllers. However, through the principle of forward invariance, CBFs can guarantee safety only when the QP is feasible and optimal [2, 3, 4]. To improve deployability, prior work has explored relaxed CBFs (R-CBF) with slack penalties [5], the interplay between input limits and barrier constraints [6, 7], higher-order CBFs to solve the high-degree dynamic system [8], and learning-driven tuning of controller and safety parameters [9]. Under environmental uncertainties, stochastic CBFs [10, 11, 12] and CVaR-based CBFs [13, 14] address tail events (like unpredictable pedestrians) at additional computational cost and conservative behaviors.
@@
Several works acknowledge that input and safety constraints can render the inner QP infeasible and pursue mitigation via parameter tuning or adaptation [9, 14]. Others address feasibility through hierarchical optimization or learning-based augmentation of feasible sets [15]. While these approaches improve feasibility, they often trade off computational budget or require substantial data collection to bias the controller toward feasible solutions, and they may not directly quantify safety when infeasibility persists under real-time limits.
This paper begins from a practical perspective, suggesting that infeasibility does not necessarily imply a loss of safety. Infeasible or suboptimal steps may arise from calculation time limitations or input constraints rather than imminent safety violations. Focusing on the safety margin with respect to the system, we propose a risk-budgeted monitor that evaluates the safety quality of the control action over a desired window length. Instead of penalizing a single suboptimal or infeasible control action, we now look over a window to certify safe behavior of the AV. This allows the system to recover to a safe state despite occasional undesired control actions that might have occurred due to computational complexity or input limitations. Inferring that monitoring the safety margin does not make the system safer, we utilize the risk monitor warning as a switch signal to determine when to use a conservative safety filter. In short, we run a relaxed CBF-QP by default for trajectory tracking. When the risk monitor detects a decline in safety quality, the filter will switch to a conservative CVaR–CBF to bound the tail risk under uncertainty. The overall architecture is shown in Fig. 1. This work makes the following key contributions.
@@
    We formalize a risk–budget monitor that provides windowed safety guarantees over a finite horizon.
@@
    We develop a monitor–driven switching mechanism between a performance–oriented CBF–QP and a conservative CVaR–CBF, enabling real-time navigation with bounded tail risk.
@@
    We theoretically and experimentally show that the risk budgeting permits controlled, transient infeasibility while preserving certified finite-horizon safety.
II Related Work
Recent work tackles the interplay between infeasibility and safety constraints from complementary angles. Ma et al. embed an exponential CBF QP as a differentiable layer so that the safety shaping is learned rather than hand-tuned [16]. A penalized slack is used only during training to keep the QP solvable while back-propagating. Manda et al. instead learn the barrier itself so that the control-invariant safe set is enlarged under actuation limits, lowering the likelihood that the safety-filter QP becomes empty online [17]; the trade-off is scalability and expressivity at the expense of limited formal robustness guarantees. Overall, these data-driven approaches address infeasibility by learning either the shaping, the tuning, or the certificate, while keeping safety as a hard constraint. However, their effectiveness can be scenario dependent and operation design domain limitation, requiring representative data and careful validation before broad deployment.
Ahmadi et al. introduce CVaR barrier functions (CVaR-BFs) for discrete-time systems that enforce safety on the distributional tail [13]. For dynamic crowded environments, Wang et al. make the feasibility–safety explicit that smaller risk levels contract the feasible set, whereas larger risk levels expand it at the cost of conservatism; they adapt the CVaR confidence level online to the minimum required for feasibility and augment with a dynamic-zone CBF that enlarges the virtual safety radius under risky relative motion, preserving the intended probabilistic guarantee [14]. Across these works, infeasibility is mitigated by learning risk parameters and exploiting local feasibility of the stepwise optimization, but universal feasibility certificates under tight actuation and aggressive interactions remain open, with computational load. [15]. In both cases, safety remains the top priority—either by fencing off infeasible states or by guaranteeing a feasible least-violation solution when perfect invariance is impossible.
The chapter of the paper is organized as follows. We first review standard and relaxed CBF formulations, then define stochastic safe sets and introduce the relaxed CVaR–CBF formulation for real-time implementation. Section IV formalizes the sliding-window risk-budgeted monitor for safety risk, and Section V details the triggering logic for mode switching. We conclude with numerical experiments and a discussion of results.
III Preliminaries
Consider a nonlinear continuous time, control-affine system given by:
@@ III Preliminaries
where $\bm{x}\in\mathcal{D}\subset\mathbb{R}^{n}$ and $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represent the system’s state and input. $\mathcal{U}$ is the set of admissible inputs. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},\quad$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are locally Lipschitz continuous.
The safe set $\mathcal{S}=\{\bm{x}:h(\bm{x})\geq 0\}$ is the zero-superlevel set of a function $h:D\subset\mathbb{R}^{n}\rightarrow\mathbb{R}$, and the system is guaranteed to be safe if the control input satisfies:
$L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}+\kappa h(\bm{x})\ \geq\ 0,$ (2)
where $\kappa$ is a user-defined positive constant that determines the minimum decay rate of $h(\bm{x})$[18]. The dynamics of the barrier function is given by:
$\dot{h}(\bm{x},\bm{u})=L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}.$ (3)
Let a safety residual be defined as $r(\bm{x},\bm{u})$:
@@
For implementing the controller in a real-time system like autonomous vehicle (AV), the continuous-time system is discretized using a zero-order hold (ZOH)
$\bm{u}(t)=u_{k},\quad\forall t\in[t_{k},t_{k+1}),$ (5)
where $u_{k}$ is the constant control input between the sampling time $t_{k}$ and $t_{k+1}$. The sampling has a uniform sampling time interval $T_{s}$ [19]. Thus, as in (5), the barrier dynamics in (3) satisfy
$\dot{h}(\bm{x},u_{k})=-\kappa h(\bm{x},u_{k})+r(\bm{x},u_{k}).$ (6)
The sample-wise lower bound is given by
$r_{k}=\inf_{t\in[t_{k},t_{k+1})}r\big(\bm{x},u_{k}\big)$ (7)
$\qquad(\text{for small }T_{s},\ r_{k}\approx r(\bm{x},u_{k})).$
Since $\bm{x}$ is a time varying function, we can rewrite (6) as:
$\dot{h}(t)\geq-\kappa h(t)+r_{k}.$ (8)
Solving this linear differential inequality yields the one-step comparison given by:
@@
where $\mu=e^{-\kappa T_{s}}\in(0,1)$ and $c=\frac{1-\mu}{\kappa}$. This is a sufficient constraint on the next-step state $h_{k+1}$ given $(h_{k},u_{k})$, ie, it is a guaranteed lower bound rather than an exact point prediction.
Remark 1
In our experiments, we choose $T_{s}$ to be sufficiently small so that the one-step bound (9) remains valid, while minimizing the effect of sampled-data mismatch.
III-A Relaxed CBF (R-CBF)
@@
The slack $\nu_{k}>0$ quantifies the instantaneous relaxation of the barrier constraint and is penalized by $\rho_{\nu}>0$. A smaller penalty admits larger (more aggressive and performance-seeking) relaxations, while a larger penalty on $\nu_{k}$ yields more conservative (safety-biased) behavior. In Section IV, we demonstrate how to utilize $\nu_{k}$ and the residual margins $r_{k}$ as supervisory signals that guide a balanced control framework between performance and safety.
III-B Stochastic Safe Sets
Given the inherent uncertainties in dynamic obstacle positions and vehicle localization, a safe set for the AV cannot be computed deterministically. To address this challenge, we transition to a stochastic formulation where safety is evaluated in terms of probabilistic risk. The obstacle state is sampled $P$ times from its distribution, denoted as $\bm{x}^{o,p}\in\mathcal{O}\subset\mathbb{R}^{n}$, where $p=1,\dots,P$. Similarly, the vehicle state is sampled $Q$ times, denoted as $\bm{x}^{q}\in\mathcal{D}\subset\mathbb{R}^{n}$ where $q=1,\dots,Q$. These independent samples are then combined to form a total of $S=P\times Q$ stochastic safe set. To represent each unique vehicle-obstacle pair, we use a double index $(p,q)$. The CBF for each safe set at time k is defined as:
$h_{k}^{i}$ $=\left\|{\mathbf{x}}_{k}^{v,q}-{\mathbf{x}}_{k}^{o,p}\right\|_{2}-D_{s},$ (11a)
where $i=1,\dots,S$ is the total number of pair $(p,q)$ and $D_{s}$ represents the minimum safe distance. The safety constraints $h_{k+1}$ and the residual $r_{k}$ can be written in stochastic formulation as:
@@ III-B Stochastic Safe Sets
$r_{k}^{i}(u_{k})$ $=L_{f}h_{k}^{i}+L_{g}h_{k}^{i}u_{k}+\kappa h_{k}^{i}.$ (11c)
@@
III-C Risk-Aware Safety Critical Control Formulation
III-C1 Probability Constraints
Due to the stochasticity in obstacle and vehicle positions, the optimization of the loss function is considered a chance-constrained optimization problem that accommodates uncertainties, as given by.
$\mathbb{P}\left(Z_{k+1}^{i}\leq 0\right)\geq\epsilon,$ (13)
where $\epsilon\in(0,1)$ represents the desired confidence level. This constraint ensures that the probability of the system remaining in a safe state (defined by a non-positive loss, $Z_{k+1}^{i}\leq 0$) is at least $\epsilon$. From (13), we can formulate the probability constraints to a standard measure of risk using Value-at-Risk (VaR). The VaR of the loss function at the confidence level $\epsilon$ can be defined as:
$\text{VaR}_{\epsilon}(Z_{k+1}^{i}):=\inf_{\gamma\in\mathbb{R}}\left\{\gamma\;\middle|\;\mathbb{P}(Z_{k+1}^{i}\leq\gamma)\geq\epsilon\right\},$ (14)
where $\gamma$ represents the candidate loss threshold being evaluated. The relationship between this candidate threshold $\gamma$ and the safety boundary is the key to this formulation. The initial probabilistic constraint in (13) establishes a specific requirement that the system must satisfy. Comparing the safety boundary to the VaR definition, $\mathbb{P}(Z_{k+1}^{i}\leq\gamma)\geq\epsilon$, it is evident that $\gamma=0$ must be one of the values that satisfy the safety condition.
This leads to the final, tractable constraint:
$\text{VaR}_{\epsilon}(Z_{k+1}^{i})\leq 0.$ (15)
This means that satisfying the constraint $\text{VaR}_{\epsilon}(Z_{k+1}^{i})\leq 0$ provides a probabilistic guarantee of safety. It ensures the system will not enter an unsafe state under the uncertainties considered with a confidence level of $\epsilon$.
However, VaR provides no information about the magnitude of the expected loss within the $(1-\epsilon)$ risk tail of the distribution. Once an outcome falls into this tail, the loss value could be close to zero or a very large number, which determines the severity of the safety violation. In turn, Conditional Value-at-Risk (CVaR) is introduced to address this limitation.
@@ III-C1 Probability Constraints
Definition 1 (Conditional Value-at-Risk). The CVaR at a confidence level $\epsilon$ is the expected loss in the $(1-\epsilon)$ tail of the distribution ($Z_{k+1}^{i}$), which occurs when the loss exceeds threshold $\text{VaR}_{\epsilon}$ [20].
$\text{CVaR}\epsilon(Z_{k+1}^{i}):=\mathbb{E}\left[Z_{k+1}^{i}\middle|Z_{k+1}^{i}\geq\text{VaR}_{\epsilon}(Z_{k+1}^{i})\right].$ (16)
From [21], the (16) can be rewritten as:
@@
$\text{CVaR}_{\epsilon}(Z_{k+1}^{i})=\inf_{\gamma\in\mathbb{R}}\left\{\gamma+\frac{1}{1-\epsilon}\,\mathbb{E}\!\left[(Z_{k+1}^{i}-\gamma)_{+}\right]\right\},$ (17)
where $(x)_{+}:=\max(0,x)$ is the positive part function. The infimum is achieved when $\gamma$ is equal to $\text{VaR}_{\epsilon}(Z_{k+1}^{i})$. The sample-based CVaR of the loss $Z_{k}^{i}$ can be found by solving for the optimal risk threshold $\gamma$.
In practice, finding a control solution that completely eliminates risk is not always feasible. Therefore, we introduce a slack variable $\nu_{k}$ to represent the accepted level of risk, and refer to the resulting method as the Relaxed CVaR–CBF. The safety filter then minimally modifies a nominal command $u_{k}^{\text{nom}}$ by solving the following quadratic programming (QP) problem:
[…]
$\text{s.t.}\quad\text{CVaR}\epsilon(Z_{k+1}^{i})\leq\nu_{k},$ $0\leq\nu_{k}\leq\overline{\nu},$
$r_{k}^{i}(u_{k})\ \geq\ -\,\nu,\quad i=1,\dots,S,$
where $\nu_{k}$ represents the upper bound of the tail severity, while $\overline{\nu}$ acts as a risk cap, limiting the maximum allowable risk for any single step. The window-level safety certificate determines the risk cap’s maximum allowable magnitude to guarantee safety over a finite time horizon, as discussed in IV-B.The coefficient $\rho_{\nu}>0$ is a penalty weight, and the term $\rho_{\nu}\nu_{k}^{2}$ in the cost function penalizes any non-zero risk. By allowing $\nu_{k}$ to be non-zero, this formulation makes the controller more robust by preventing infeasibility in challenging situations, while $\nu_{k}$ provides a direct, quantifiable measure of the risk taken in that step.
IV Risk-Budget Monitor
As the optimization in (18) allows for non-zero risk, a balance can be struck between performance, safety, and the prevention of solution infeasibility. Furthermore, a single infeasible step or a momentary non-zero risk might not immediately lead the AV into an unsafe area, especially when the system operates with a safety buffer. To manage this trade-off effectively, this section defines the criteria for bad step detection and the mechanism of the risk window monitor.
IV-A Bad Step detection
Before defining the bad step condition, we must first define the parameter $\delta$, which represents a pre-defined safety margin required for a good step. and it should be a positive number $\delta>0$. A step is considered good only if the residual $r_{k}$ is greater than or equal to $\delta$, effectively depositing a guaranteed safety allowance into the system.
Accordingly, a step is classified as a bad step if it meets either of the following two conditions:
  • Condition I: $\nu_{k}>\bar{\nu}$
    The variable $\nu_{k}$ quantifies the controller’s acceptance of finding a feasible solution, and $\bar{\nu}$ is the worst allowance risk. If $\nu_{k}$ is greater than $\bar{\nu}$, it signifies that the chosen control action has exceeded the maximum allowable risk. Therefore, the step is classified as a bad step.
  • Condition II: $r_{k}<\delta$

### v2

# Risk-Budgeted Control Framework for Improved Performance and Safety in Autonomous Vehicles

@@ Abstract
This paper presents a hybrid control framework with a risk-budgeted monitor for safety-certified autonomous driving. A sliding-window monitor tracks insufficient barrier residuals and triggers switching from a relaxed control barrier function (R-CBF) to a more conservative conditional value-at-risk CBF (CVaR-CBF) when the safety margin deteriorates. Two real-time triggers are considered: feasibility-triggered (FT), which activates CVaR-CBF when the R-CBF problem is reported infeasible, and quality-triggered (QT), which switches when the residual falls below a prescribed safety margin.
The framework is evaluated with model predictive control (MPC) under vehicle localization noise and obstacle position uncertainty across multiple AV-pedestrian interaction scenarios with 1,500 Monte Carlo runs. In the most challenging case with 5 m pedestrian detection uncertainty, the proposed method achieves a 94–96% collision-free success rate over 300 trials while maintaining the lowest mean cross-track error (CTE = 3.2–3.6 m), indicating faster trajectory recovery after obstacle avoidance and a favorable balance between safety and performance.
@@ I Introduction
Safety under uncertainty is a fundamental requirement in autonomous vehicle (AV) control, and the challenge is amplified when safety filters are implemented as real-time optimization routines operating at high frequencies (e.g., 20205050 Hz) [10]. Under strict computational budgets, the online solver may return (i) a feasible (near-)optimal solution, (ii) a feasible but non-optimal iterate due to early termination or imperfect convergence, or (iii) a solver-reported infeasibility outcome. In such real-time settings, reliably quantifying the degree of suboptimality—or determining whether a returned iterate remains sufficiently safe—on a step-by-step basis is generally nontrivial. Accordingly, we study safety under non-ideal solver outcomes, encompassing both suboptimal feasible iterates and solver-reported infeasibility.
Control Barrier Functions (CBFs) formulated as quadratic programs (QPs) provide a principled mechanism for enforcing forward invariance, guaranteeing safety when the CBF-QP remains feasible and is solved with sufficient optimality [5, 6, 8]. To improve deployability under input constraints and modeling complexity, prior work has considered relaxed CBFs with slack penalties [17], input-aware formulations [9, 1], higher-order CBFs [14], and learning-based tuning of controller or safety parameters [12, 18]. While these methods mitigate conservatism and feasibility issues in deterministic settings, environmental uncertainty can still erode safety margins or induce overly conservative behavior.
@@
Under uncertainty, stochastic and CVaR-based CBF formulations explicitly address distributional tail risk [7, 2, 16]. These approaches regulate confidence levels to trade off feasibility and conservatism, sometimes adapting risk online to preserve probabilistic safety guarantees. However, chance-constrained and stochastic safety filters may still be overly conservative in practice and often incur higher computational cost. Although CVaR–CBF recovers the deterministic CBF case in the limit, enforcing risk-aware optimization at every step is often unnecessary for high-rate real-time deployment.
Related efforts include learning-based feasibility recovery and probabilistic barrier methods that adapt constraints or risk parameters online (e.g., [2]), as well as studies that provide safety guarantees under solver-induced non-optimality (e.g., early termination or imperfect convergence) [3, 15]. In contrast to per-step feasibility approaches, we certify safety over a finite horizon using window-level risk budgeting under occasional non-ideal solver outcomes.
This paper adopts a practical perspective: non-ideal solver outcomes do not necessarily imply immediate safety loss, but repeated occurrences can gradually erode the safety margin. Instead of evaluating safety at each step, we monitor the evolution of the safety margin over a sliding window. The proposed risk-budgeted monitor evaluates safety quality over a finite horizon, allowing occasional undesired control actions due to computational or input limitations while still certifying windowed safety.
Since monitoring alone cannot enforce safety, the monitor output is used as a switching signal in a hybrid architecture that localizes conservatism to periods of accumulated risk. The nominal mode employs a lightweight relaxed CBF-QP to preserve tracking performance, while a conservative CVaR–CBF mode is activated when the monitored safety quality degrades. The overall architecture is illustrated in Fig. 1.
The main contributions of this work are:
@@
    A finite-horizon risk-budget formulation that certifies safety under transient non-optimal solutions.
@@
    A monitor-driven switching mechanism between performance-oriented CBF-QP and conservative CVaR–CBF filters.
@@
    Theoretical analysis and simulations demonstrating balanced safety and performance under uncertainty.
II Preliminaries
Consider a nonlinear control–affine system
@@ II Preliminaries
where $\bm{x}\in\mathcal{D}\subset\mathbb{R}^{n}$ and $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ denote the state and input, respectively. Let the safe set be defined as
$\mathcal{S}=\{\bm{x}:h(\bm{x})\geq 0\},$
where $h:\mathcal{D}\rightarrow\mathbb{R}$ is a continuously differentiable function. Safety is guaranteed if the control input satisfies the CBF condition [4]
$L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}+\kappa h(\bm{x})\geq 0.$ (2)
Define the safety residual
@@
For real–time implementation, the controller is applied with zero–order hold, $\bm{u}(t)=u_{k}$ for $t\in[t_{k},t_{k+1})$. Under this sampling scheme, the barrier dynamics satisfy
$\dot{h}(t)\geq-\kappa h(t)+r_{k},$ (4)
where $r_{k}=\inf_{t\in[t_{k},t_{k+1})}r(\bm{x},u_{k})$.
Solving (4) yields the one–step comparison inequality
@@
where $\mu=e^{-\kappa T_{s}}\in(0,1)$ and $c=(1-\mu)/\kappa$.
II-A Relaxed CBF (R-CBF)
@@
The slack $\nu_{k}>0$ quantifies the instantaneous relaxation of the barrier constraint and is penalized by $\rho_{\nu}>0$. A smaller penalty admits larger (more aggressive and performance-seeking) relaxations, while a larger penalty on $\nu_{k}$ yields more conservative (safety-biased) behavior. In Section III, we demonstrate how to utilize $\nu_{k}$ and the residual margins $r_{k}$ as supervisory signals that guide a balanced control framework between performance and safety.
II-B Stochastic Safe Sets
Given uncertainty in obstacle and vehicle states, safety is evaluated in a stochastic setting. Let $\bm{x}^{o,p}\in\mathcal{O}$ and $\bm{x}^{v,q}\in\mathcal{D}$ denote $P$ obstacle samples and $Q$ vehicle samples, respectively, yielding $S=PQ$ vehicle–obstacle pairs indexed by $i$. The CBF for each pair at time $k$ is
$h_{k}^{i}$ $=\left\|{\bm{x}}_{k}^{v,q}-{\bm{x}}_{k}^{o,p}\right\|_{2}-D_{s},$ (7a)
@@ II-B Stochastic Safe Sets
$r_{k}^{i}(u_{k})$ $=L_{f}h_{k}^{i}+L_{g}h_{k}^{i}u_{k}+\kappa h_{k}^{i},$ (7c)
where $i=1,\dots,S$ and $D_{s}$ is the minimum safe distance.
@@
II-C Risk-Aware Safety Critical Control Formulation
II-C1 Probability Constraints
Due to uncertainty in obstacle and vehicle positions, safety is imposed through the chance constraint
$\mathbb{P}\left(Z_{k+1}\leq 0\right)\geq\epsilon,$ (9)
where $\epsilon\in(0,1)$ is the confidence level. Let $Z_{k+1}$ denote the random loss at the next step, with sampled realizations $\{Z_{k+1}^{i}\}_{i=1}^{S}$. The chance constraint in (9) can be expressed using the Value-at-Risk (VaR) measure.
$\text{VaR}_{\epsilon}(Z_{k+1}):=\inf_{\gamma\in\mathbb{R}}\left\{\gamma\;\middle|\;\mathbb{P}(Z_{k+1}\leq\gamma)\geq\epsilon\right\},$ (10)
where $\gamma$ represents the candidate loss threshold being evaluated. Thus, the chance constraint is equivalent to requiring that the $\epsilon$-quantile of the loss distribution does not exceed zero. Hence, the probabilistic safety requirement is
$\text{VaR}_{\epsilon}(Z_{k+1})\leq 0.$ (11)
While VaR specifies the $\epsilon$-quantile of the loss distribution, it does not account for the severity of losses in the $(1-\epsilon)$ tail. To capture this tail risk, we adopt Conditional Value-at-Risk (CVaR).
@@ II-C1 Probability Constraints
The Conditional Value-at-Risk (CVaR) at confidence level $\epsilon$ is the expected loss in the $(1-\epsilon)$ tail [13]:
$\text{CVaR}_{\epsilon}(Z_{k+1}):=\mathbb{E}\!\left[Z_{k+1}\mid Z_{k+1}\geq\text{VaR}_{\epsilon}(Z_{k+1})\right].$ (12)
Using the Rockafellar–Uryasev representation [11],
@@
$\text{CVaR}_{\epsilon}(Z_{k+1})=\inf_{\gamma\in\mathbb{R}}\left\{\gamma+\frac{1}{1-\epsilon}\mathbb{E}\!\left[(Z_{k+1}-\gamma)_{+}\right]\right\},$ (13)
where $(x)_{+}:=\max(0,x)$.
In practice, enforcing zero risk at every step may render the safety filter infeasible under uncertainty. We therefore introduce a nonnegative slack variable $\nu_{k}$ to quantify the admitted risk at time $k$, and define the resulting filter as the Relaxed CVaR–CBF. Given a nominal command $u_{k}^{\mathrm{nom}}$, the control input is obtained by solving
$\min_{u_{k}\in\mathcal{U},\ \nu_{k}\in\mathbb{R}}$ $\frac{1}{2}\|u_{k}-u_{k}^{\mathrm{nom}}\|^{2}+\rho_{\nu}\nu_{k}^{2},$ (14)
s.t. $\mathrm{CVaR}_{\epsilon}\!\left(Z_{k+1}\right)\leq\nu_{k},$
$r_{k}^{i}(u_{k})\geq-\,\nu_{k},\quad i=1,\dots,S,$
$0\leq\nu_{k}\leq\overline{\nu},$
where $\rho_{\nu}>0$ penalizes admitted risk. The constraint $\mathrm{CVaR}_{\epsilon}(Z_{k+1})\leq\nu_{k}$ bounds the tail risk, while $r_{k}^{i}(u_{k})\geq-\nu_{k}$ enforces a uniform per-scenario lower bound on the residual required by the sliding-window analysis in Section III. The upper bound $\overline{\nu}$ specifies the stepwise risk cap, whose admissible magnitude is determined by the window-level safety certificate in Section III-B.
Under the sample-based Rockafellar–Uryasev epigraph reformulation, (14) becomes a convex quadratic program.
III Risk-Budget Monitor
As the optimization in (14) allows non-zero risk through relaxation, a trade-off can be achieved between performance and safety. However, under real-time computational constraints, the solver may occasionally return a non-optimal solution, such as a suboptimal convergence or a temporarily elevated risk level. Importantly, a single non-optimal step or transient increase in risk does not necessarily drive the AV into an unsafe region, particularly when a safety margin is maintained.
To manage this trade-off systematically, this section formalizes the notion of a bad step and introduces a sliding-window risk monitor to evaluate safety quality over a finite horizon.
III-A Bad Step detection
We first define $\delta>0$ as the desired per-step safety margin. A step is considered good if the residual satisfies $r_{k}\geq\delta$, which effectively deposits a safety allowance into the system. We further assume the safety filter enforces a uniform residual lower bound $r_{k}\geq-\overline{\nu}$ at every step (14). A step is classified as a bad step if it meets the following condition:
  • Condition: $r_{k}<\delta$